OpenStudy (anonymous):

ELLIPSE/HYPERBOLA: I'm so confused please help Is this an ellipse or a hyperbola and is it horizontal or vertical

OpenStudy (anonymous):

\[\frac{ x^2 }{16}-\frac{ y^2 }{ 9}=1\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Graph the points, by first solving for y algebraically.

OpenStudy (anonymous):

I dont know why i have to solve for any variable?

OpenStudy (dan815):

ellipse is x^2/a + y^2/b=1

OpenStudy (anonymous):

you don't have to solve for any variable, it just gives you an idea when graphing what it looks like.

OpenStudy (anonymous):

@dan815 so it has to be a hyperbola? but the greater value is under the x?

OpenStudy (dan815):

ya its a hyperbola

OpenStudy (dan815):

change the x and y from values if u want to see more of the graph

OpenStudy (dan815):

when y is 0, x values must be the same for hyperbola and ellipse

OpenStudy (dan815):

but when x = 0 u will get imaginary solution for y for a hyperbola

OpenStudy (anonymous):

thanks you

OpenStudy (dan815):